OpenStudy (anonymous):

Find the x-intercepts of the parabola with vertex (-3,16) and y-intercepts (0,-20). Write your answer in this form: (x1,y1), (x2,y2) If necessary, round to the nearest hundredth.

OpenStudy (loser66):

What is the vertex form of a parabola?

OpenStudy (anonymous):

(y-k)=a(x-h)^2

OpenStudy (loser66):

ok, replace h, k what do you have?

OpenStudy (anonymous):

-3?

OpenStudy (welshfella):

h = -3 and what about k?

OpenStudy (anonymous):

k would be 16. right?

OpenStudy (loser66):

yup

OpenStudy (welshfella):

so we have y -16 = a(x + 3)^2

OpenStudy (anonymous):

okay so what now?

OpenStudy (welshfella):

you know have to find the value of a you know that when x = 0 , y = -20 ( the y-intercept) so how do you think we can find the value of a using these values of x and y ?

OpenStudy (welshfella):

* you now

OpenStudy (welshfella):

hmm. another one disappeared..